Verify The Equation is an Identity
cosx*cos(-x)-sinx*sin(-x)=1
how do I do it???

Answer by lwsshak3(11628)

Verify The Equation is an Identity
cosx*cos(-x)-sinx*sin(-x)=1
..
cos(-x)=cos(x) (cos is an even function)
sin (-x)=-sinx (sin is an odd function)
..
cosx*cos(-x)-sinx*sin(-x)
cosx*cosx-(sinx*(-sinx)=cos^2x+sin^2x=1